mikewwe13:

Justin Moore has scores of 72, 67, 82, and 79 on his algebra tests. A. Use an inequality to find the scores he must make on the final exam to pass the course with an average of 77 or higher, given that the final exam counts as two tests. B. Explain the meaning of the answer to part a

jhonyy9:

so i think \[\frac{ \left( 72 +67 +82 +79 +a +b \right) }{ 6 } >= 77\] solve this inequality for (a +b) what represent these two tests of the final examen

jhonyy9:

@Laylalyssa any idea ? ty

mikewwe13:

(72 + 67 + 82 + 79 + x + x ) / 6 = 77

mikewwe13:

i have to graph it, i'm doing inequality.

mikewwe13:

@jhonyy9

jhonyy9:

why x ,x ? do you wan calcule them like 2 the same ,equal results on the final examen ,on these two tests ?